# Fan distributions via Tverberg partitions and Gale duality

### v2

@@ 4. Distributions by a Single Fan
To that end, we first show that any $r-1$ of $\beta_{1},\ldots,\beta_{r}$ are linearly dependent in $\mathbb{R}^{n-d-1}$. In particular, each $\beta_{j}$ is non-zero, so each $H^{\prime\prime}_{j}$ is a hyperplane in $\mathbb{R}^{n-d-1}$. Without loss of generality, consider $\beta_{1},\ldots,\beta_{r-1}$ and suppose that $\sum_{j=1}^{r-1}\mu_{j}\beta_{j}=0$. We have $\langle\sum_{j=1}^{r-1}\mu_{j}\alpha_{j},g_{n+1}\rangle=0$ because $g_{n+1}\in H_{j}$ for all $j\in[r]$. As $g_{n+1}=(-x,-n)$ where $x=\sum_{i=1}^{n}x_{i}\in\mathbb{R}^{n-d-1}$, we have $\langle\sum_{j-1}^{r}\mu_{j}\alpha_{j},g_{n+1}\rangle=-n\sum_{j=1}^{r-1}\mu_{j}\gamma_{j}$. Thus $\sum_{j=1}^{r-1}\mu_{j}\gamma_{j}=0$, and since $\sum_{j=1}^{r-1}\mu_{j}\beta_{j}=0$ by assumption we have that $\sum_{j=1}^{r-1}\mu_{j}\alpha_{j}=0$ as well. However, any $r-1$ of the $\alpha_{j}$ are linearly independent in $\mathbb{R}^{n-d}$, so each $\mu_{j}$ is zero and $\beta_{1},\ldots,\beta_{r-1}$ are linearly independent in $\mathbb{R}^{n-d-1}$. On the other hand, $\beta_{1},\ldots,\beta_{r}$ are linearly dependent because $\alpha_{1},\ldots,\alpha_{r}$ are. Finally, we show that $\cap_{j=1}^{r}H^{\prime}_{j}\neq\emptyset$ and so that $H_{j}^{\prime\prime}\neq\emptyset$ as well. But $g_{n+1}=(-x,-n)\in\cap_{j=1}^{r}H_{j}$, so $(\frac{-x}{n},1)\in\cap_{j=1}^{r}H_{j}^{\prime}\neq\emptyset$.
@@
In the complex setting, we have a complex regular $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{C}^{n-d}$ centered about a linear complex hyperplane $H_{\mathbb{C}}$. As above, we let $H_{\mathbb{C}}^{\prime}=H_{\mathbb{C}}\cap(\mathbb{C}^{n-d-1}\times\{1\})$, $B_{j}^{\prime}=B_{j}\cap(\mathbb{C}^{n-d-1}\times\{1\})$ for each $j\in[r]$, and we let $H^{\prime\prime}_{\mathbb{C}}$ and $B^{\prime\prime}_{j}$ be their corresponding projections onto $\mathbb{C}^{n-d-1}$. We claim that $F_{r}^{\prime\prime}=\cup_{j=1}^{r}B^{\prime\prime}_{j}$ is a regular $r$-fan in $\mathbb{C}^{n-d-1}$ centered about the complex hyperplane $H^{\prime\prime}_{\mathbb{C}}$. To verify this, we have $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$, where $\alpha=(\beta,\gamma)\in\mathbb{C}^{n-d-1}\times\mathbb{C}$ is non-zero. Thus $H^{\prime\prime}_{\mathbb{C}}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle=-\gamma\}$ and $B^{\prime}_{j}=\{(z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle+\gamma=t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. Thus $H_{\mathbb{C}}^{\prime\prime}$ and $F_{r}^{\prime\prime}$ are as claimed provided $\beta\neq 0$. But $g_{n+1}=(-x,-n)\in H_{\mathbb{C}}$ where $x=\sum_{i=1}^{n}g_{i}$. Thus $\langle(x,-n),(\beta,\gamma)\rangle=0$ and so $\langle\beta,x\rangle=n\gamma$. But if $\beta=0$, then $\gamma=0$ and therefore $\alpha=0$ as well. Thus $F^{\prime\prime}_{r}$ is indeed a regular $r$-fan centered about the complex hyperplane $H_{\mathbb{C}}^{\prime\prime}$.
@@
To complete the proof, we again let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$ and show that the interior $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$ of any $B^{\prime\prime}_{j}$ does not contain any element of $\mathcal{F}$. So let $F\in\mathcal{F}$ and let $F^{\prime}$ be the corresponding element of $\mathcal{F}^{\prime}$. We have that $|F^{\prime}|=|F|$ and $|F^{\prime}\cap I_{j}|=|F\cap\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})|$ for all $j\in[r]$. Since $F^{\prime}$ is not contained in any of the $I_{j}$, we conclude that $F$ is not contained in any of the $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$. ∎
@@
We conclude this section by proving our colorful distribution result. This follows the same procedure as before, except we substitute Sarkaria’s theorem with the Optimal Colored Tverberg theorem of  [8] below.
@@
Let $d\geq 1$ be an integer, let $r\geq 2$ be an integer such that $r+1$ is prime, and let $n\geq r(d+1)$. If $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and $C_{1}\sqcup\cdots\sqcup C_{d+1}$ is a $(d+1)$-coloring of the vertex set of $\Delta_{n}$ such that each $|C_{i}|\geq r$ for all $i\in[d+1]$, then there exists a Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that $|\sigma_{j}\cap C_{i}|\leq 1$ for all $j\in[r]$ and all $i\in[d+1]$.
@@
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. We let $n\geq r(d+1)$ if $\mathbb{K}=\mathbb{R}$ and $n\geq r(2d+1)$ if $\mathbb{K}=\mathbb{C}$ and we suppose that $X=\{x_{1},\ldots,x_{n}\}$ is a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$. In the real case we have that $X=X_{1}\sqcup\cdots\sqcup X_{d+1}$ is a $(d+1)$-coloring with $|X_{k}|\geq r$ for all $k\in[d+1]$, and likewise in the complex case that with $d+1$ replaced by $12d+$ 1. As in the proof of Theorem 1.4, lift $X$ to the hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ of height $1$ in $\mathbb{K}^{n-d}$. As there, the points $g_{1}=(x_{1},1)\,\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$ are the Gale transform of some sequence $a_{1},\ldots,a_{n+1}$ of affinely spanning points in $\mathbb{K}^{d}$. As before, we let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the (real) affine linear map defined by $f(i)=a_{i}$ for all $i\in[n+1]$. If $\mathbb{K}=\mathbb{R}$, we define $C_{k}=\{i\in[n]\mid x_{i}\in X_{k}\}$ for each $k\in[d+1]$. Thus $C_{1}\sqcup$ is a $(d+1)$-coloring of the vertex set of $\Delta_{n-1}$ into $d+1$ color classes of size at least $r$ each. We proceed analogously in the complex setting. Applying Theorem 4.1 to the restriction of $f$ to $\Delta_{n-1}$, we have a proper $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f\mid_{\Delta_{n-1}}$ such that the vertex set $I_{j}\subset[n]$ of each $\sigma_{j}$ contains at most one element of each color class.

### v3

@@ 4. Distributions by a Single Fan
Now let $\mathbb{K}=\mathbb{R}$. For each $j\in[r]$ we have $H_{j}=\langle\alpha_{j}\rangle^{\perp}$ where $\alpha_{j}=(\beta_{j},\gamma_{j})\in\mathbb{R}^{n-d-1}\times\mathbb{R}$ is non-zero. The linear conical $r$-fan is of the form $F_{r}=\cup_{j=1}^{r}B_{j}$, where $B_{j}=H_{j}^{+}\cap A_{j}$, $A_{j}=\cap_{i\neq j,j-1}H_{i}$, and $H_{j}^{+}$ is some half-space determined by $H_{j}$. For each $j\in[r]$, let $H^{\prime}_{j}=H_{j}\cap(\mathbb{R}^{n-d-1}\times\{1\})$ be the intersection of $H_{j}$ with the plane of height one in $\mathbb{R}^{n-d}$. Thus $H^{\prime}_{j}=\{(x,1)\in\mathbb{R}^{n-d-1}\times\{1\}\mid\langle\beta_{j},x\rangle=-\gamma_{j}\}$. Likewise, let $B^{\prime}_{j}=B_{j}\cap(\mathbb{R}^{n-d-1}\times\{1\})$ for all $j\in[r]$. Letting $H_{j}^{\prime\prime}$ and $B_{j}^{\prime\prime}$ be the corresponding projections of $H_{j}^{\prime}$ and $B_{j}^{\prime}$ onto $\mathbb{R}^{n-d-1}$, we will show that $F^{\prime\prime}_{r}=\cup_{j=1}^{r}B_{j}^{\prime\prime}$ is a conical $r$-fan in $\mathbb{R}^{n-d-1}$. To that end, we first show that any r1r-1 of $\beta_{1},\ldots,\beta_{r}$ are linearly dependent in $\mathbb{R}^{n-d-1}$. In particular, each $\beta_{j}$ is non-zero, so each $H^{\prime\prime}_{j}$ is an affine hyperplane in $\mathbb{R}^{n-d-1}$. Without loss of generality, consider $\beta_{1},\ldots,\beta_{r-1}$ and suppose that $\sum_{j=1}^{r-1}\mu_{j}\beta_{j}=0$. We have $\langle\sum_{j=1}^{r-1}\mu_{j}\alpha_{j},g_{n+1}\rangle=0$ because $g_{n+1}\in H_{j}$ for all $j\in[r]$. As $g_{n+1}=(-x,-n)$ where $x=\sum_{i=1}^{n}x_{i}\in\mathbb{R}^{n-d-1}$, we have $\langle\sum_{j-1}^{r}\mu_{j}\alpha_{j},g_{n+1}\rangle=-n\sum_{j=1}^{r-1}\mu_{j}\gamma_{j}$. Thus $\sum_{j=1}^{r-1}\mu_{j}\gamma_{j}=0$, and since $\sum_{j=1}^{r-1}\mu_{j}\beta_{j}=0$ by assumption we have that $\sum_{j=1}^{r-1}\mu_{j}\alpha_{j}=0$ as well. However, any r1r-1 of the $\alpha_{j}$ are linearly independent in $\mathbb{R}^{n-d}$, so each $\mu_{j}$ is zero and $\beta_{1},\ldots,\beta_{r-1}$ are linearly independent. On the other hand, $\beta_{1},\ldots,\beta_{r}$ are linearly dependent because $\alpha_{1},\ldots,\alpha_{r}$ are. Finally, we show that $\cap_{j=1}^{r}H^{\prime}_{j}\neq\emptyset$ and so that $H_{j}^{\prime\prime}\neq\emptyset$ as well. But $g_{n+1}=(-x,-n)\in\cap_{j=1}^{r}H_{j}$, so $(\frac{x}{n},1)\in\cap_{j=1}^{r}H_{j}^{\prime}\neq\emptyset$.
@@
In the complex setting, we have a complex regular $r$-fan $F_{r}=\cup_{j=1}^{r}B_{j}$ in $\mathbb{C}^{n-d}$ centered about a linear complex hyperplane $H_{\mathbb{C}}$. As above, we let $H_{\mathbb{C}}^{\prime}=H_{\mathbb{C}}\cap(\mathbb{C}^{n-d-1}\times\{1\})$, $B_{j}^{\prime}=B_{j}\cap(\mathbb{C}^{n-d-1}\times\{1\})$ for each $j\in[r]$, and we let $H^{\prime\prime}_{\mathbb{C}}$ and $B^{\prime\prime}_{j}$ be their corresponding projections onto $\mathbb{C}^{n-d-1}$. We claim that $F_{r}^{\prime\prime}=\cup_{j=1}^{r}B^{\prime\prime}_{j}$ is a regular $r$-fan in $\mathbb{C}^{n-d-1}$ centered about the complex affine hyperplane $H^{\prime\prime}_{\mathbb{C}}$. To verify this, we have $H_{\mathbb{C}}=\langle\alpha\rangle^{\perp}$, where $\alpha=(\beta,\gamma)\in\mathbb{C}^{n-d-1}\times\mathbb{C}$ is non-zero. Thus $H^{\prime\prime}_{\mathbb{C}}=\{z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle=-\gamma\}$ and $B^{\prime\prime}_{j}=\{(z\in\mathbb{C}^{n-d-1}\mid\langle\beta,z\rangle=-\gamma+t\omega_{r}^{j}\,\,\text{for some}\,\,t\geq 0\}$. We therefore have that $H_{\mathbb{C}}^{\prime\prime}$ and $F_{r}^{\prime\prime}$ are as claimed provided $\beta\neq 0$. But $g_{n+1}=(-x,-n)\in H_{\mathbb{C}}$, where $x=\sum_{i=1}^{n}x_{i}$, and so $\langle\beta,x\rangle=-n\gamma$. If $\beta=0$, then $\gamma=0$ and therefore $\alpha=0$ as well, a contradiction. Thus $\beta\neq 0$ and $F^{\prime\prime}_{r}$ is indeed a complex regular $r$-fan with center $H_{\mathbb{C}}^{\prime\prime}$.
@@
Again let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. Since $g_{1},\ldots,g_{n+1}$ are distributed according to the $I_{j}$, for each $i\in[n]$ we have (1) for each $j\in[r]$ that $x_{i}$ lies in $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$ if $i\in I_{j}$ and (2) that $x_{i}$ lies in the center of $F_{r}^{\prime\prime}$ for all $i\notin\sqcup_{j=1}^{r}I_{j}$. In particular, $F^{\prime\prime}_{r}$ distributes $X$. To complete the proof, we show that no member of $\mathcal{F}$ is contained in any $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$. So let $j\in[r]$, let $A\in\mathcal{F}$, and let $A^{\prime}$ be the corresponding element of $\mathcal{F}^{\prime}$. As $A^{\prime}$ is not contained in $I_{j}$, there exists some $i\in[n]$ such that $i\in A^{\prime}\setminus I_{j}$. Thus $x_{i}\in A\setminus\operatorname{\mathrm{Int}}(B_{j}^{\prime\prime})$ and so $A$ is not contained in $\operatorname{\mathrm{Int}}(B^{\prime\prime}_{j})$. ∎
@@
We conclude this section by proving our rainbow distribution result. This follows the same procedure as before, except now we substitute Sarkaria’s theorem with the following result [9, Corollary 2.3] which gives optimal results to the “colored Tverberg” problem of Bárány and Larman [2].
@@
Let $d\geq 1$ be an integer, let $r\geq 2$ be an integer such that $r+1$ is prime, and let $n\geq r(d+1)$. If $f\colon\Delta_{n-1}\rightarrow\mathbb{R}^{d}$ is a continuous map and $C_{1}\sqcup\cdots\sqcup C_{d+1}$ is a $(d+1)$-coloring of the vertex set of $\Delta_{n}$ such that $|C_{i}|\geq r$ for all $i\in[d+1]$, then there exists a Tverberg $r$-tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f$ such that $|\operatorname{\mathrm{Vert}}(\sigma_{j})\cap C_{i}|\leq 1$ for all $j\in[r]$ and all $i\in[d+1]$.
@@
Let $\mathbb{K}=\mathbb{R}$ or $\mathbb{C}$. We let $n\geq r(d+1)$ if $\mathbb{K}=\mathbb{R}$ and $n\geq r(2d+1)$ if $\mathbb{K}=\mathbb{C}$, and we suppose that $X=\{x_{1},\ldots,x_{n}\}$ is a set of $n$ affinely spanning points in $\mathbb{K}^{n-d-1}$. In the real case we have that $X=X_{1}\sqcup\cdots\sqcup X_{d+1}$ is a $(d+1)$-coloring with $|X_{k}|\geq r$ for all $k\in[d+1]$, and likewise in the complex case with $d+1$ replaced by $2d+1$. As in the proof of Theorem 1.4, lift $X$ to the $\mathbb{K}$-affine hyperplane $\mathbb{K}^{n-d-1}\times\{1\}$ in $\mathbb{K}^{n-d}$. As there, the points $g_{1}=(x_{1},1)\,\ldots,g_{n}=(x_{n},1),g_{n+1}=-\sum_{i=1}^{n}g_{i}$ are the Gale transform of some sequence $a_{1},\ldots,a_{n+1}$ of affinely spanning points in $\mathbb{K}^{d}$. As before, we let $f\colon\Delta_{n}\rightarrow\mathbb{K}^{d}$ be the (real) affine linear map defined by $f(i)=a_{i}$ for all $i\in[n+1]$. If $\mathbb{K}=\mathbb{R}$, we define $C_{k}=\{i\in[n]\mid x_{i}\in X_{k}\}$ for each $k\in[d+1]$. Thus $C_{1}\sqcup\cdots\sqcup C_{d+1}$ is a $(d+1)$-coloring of the vertex set of $\Delta_{n-1}$ into $d+1$ color classes, each of which has size at least $r$. We proceed analogously in the complex setting. Applying Theorem 4.1 to the restriction of $f$ to $\Delta_{n-1}$, we have a proper $r$-Tverberg tuple $(\sigma_{1},\ldots,\sigma_{r})$ for $f\mid_{\Delta_{n-1}}$ such that the vertex set $I_{j}\subset[n]$ of each $\sigma_{j}$ contains at most one element of each color class.